OpenStudy (anonymous):

i have a problem. Find the volume of the solid obtained by rotating the region bounded by the curves. y=x^2 and y=1 about the line y=3

OpenStudy (anonymous):

i used the washer method getting two radius (3-x^2) and (3-1)=2 \[\Pi \int\limits_{-1}^{1} (3-x ^{2})^{2}-(2)^{2}dx\] i started of with this

OpenStudy (mathmate):

y=x^2 and y=1 intersect at (0,0) and (1,1). If the area is rotated about y=3, that would make many washers, therefore the washer method. Take

OpenStudy (mathmate):

Sorry, I though it was y=x, but it is y=1, so the limits are indeed from -1 to +1.

OpenStudy (anonymous):

yea after i got the interval i factored out the two and got (x^4-6x^2+9) - (4) dx

OpenStudy (anonymous):

then i took the integral of x^4-6x^2+5 and it came out to be x^5/5-6x^3/3+5x

OpenStudy (anonymous):

after that i plugged in 1 and -1 and the answer i got was wrong

OpenStudy (mathmate):

I don't get the same expression as you. It's (x^5-10x^3+25x)/10. The numeric answer is 16/5.

OpenStudy (mathmate):

multiplied by two pi to get 32pi/5.

OpenStudy (anonymous):

i used washer method and got it wrong bt you used shell method...idk how to use shell for this one

OpenStudy (mathmate):

I used the washer method also. The integral is: \[\int\limits_{-1}^{1}2\pi((1-x^2)(3-(1+x^2)/2)dx\]

OpenStudy (anonymous):

whatttt?!?!! we never learned washer method with 2pi before

OpenStudy (mathmate):

I suppose the 2pi has to come in whenever you rotate something, right?

OpenStudy (anonymous):

yea whenever its been rotated bt we were given a formula like this pi(outer radius)^2 - pi(inner radius)^2 dx

OpenStudy (mathmate):

It actually comes to the same thing. You would probably put: \[\pi \int\limits_{-1}^{1}((3-x^2)^2-2^2)dx\] and it gives the same answer of 32pi/5.

OpenStudy (anonymous):

ohhhh ok thanks for the helpppp!!

OpenStudy (mathmate):

You're welcome!